Use the following two equation, Using the two equations to find the values of Qd, Qs, the market situation (Shortage/Surplus/Equilibrium), and the Value of shortage or surplus if any, at the following prices: 3.044, 3.65, 3.88, and 3.95.  


QS = 972 + 103.5P 


 QD = 1722 – 141.5P 




1
Expert's answer
2022-05-17T17:22:10-0400

"Q_s=972+103.5P,"

"Q_d=1722-141.5P,"

"Q_s=Q_d,"

"972+103.5P=1722-141.5P,"

"750=245P,"

"P=3.06,"

"Q=1288.8."